OpenStudy (anonymous):

(-2x^2+5x-8)+(-9-7x+8x^2)+(3+7x^2-2x)

OpenStudy (anonymous):

remove parentheses and combine like terms

OpenStudy (anonymous):

\[-2x^2+5x-8-9-7x+8x^2+3+7x^2-2x\] \[8x^2+7x^2-2x^2=13x^2\] \[5x-7x-2x=-4x\] \[-8-9+3=-14\] "final answer" \[13x^2-4x-14\]